A 5kg object is hanging by a 1.5m wire when it is suddenly hit by a 3kg missile traveling horizontally at 12m/s. The missile embeds itself in the object during the collision. What is the tension in the rod immediately after the collision?


1
Expert's answer
2021-12-14T13:33:05-0500

"mv=mv'+MV"


"mv^2\/2=mv'^2\/2+MV^2\/2" . So, have


"3\cdot12=3v'+5V"


"3\cdot12^2=3v'^2+5V^2\to" "V=5.14\ (m/s)"


"F=ma\to F=m\cdot\frac{V^2}{r}=5\cdot\frac{5.14^2}{1.5}=88\ (N)"


"T=F+mg=88+5\cdot9.8=137\ (N)" . Answer